The working mechanism of a pendulum clock is based on the principle of Seconds
Pendulum in keeping the time right.
(I) What is the length of the clock’s pendulum in order to keep the time accurate
on Earth at sea level?
(II)Given that, the length of a pendulum’s clock is 1.50 m however it is running
slow by 30 s in each day. How much you should lengthen or shorten it? Why?
1
Expert's answer
2020-12-06T17:28:15-0500

i)


"T=2\pi\sqrt{\frac{l}{g}}\\2=2\pi\sqrt{\frac{l}{9.8}}\\l=0.99\ m"

ii) We need to shorten it by


"d=1.50-0.99=0.51\ m"